OpenStudy (anonymous):

What is the length of the hypotenuse of a right triangle whose legs have lengths of 5 and 12?

OpenStudy (mathstudent55):

This is a problem for the Pythagorean theorem. \(a^2 + b^2 = c^2\) where a = 5 and b = 12.